# CRAG - Comprehensive RAG Benchmark

### v1

@@ 5.2 State-of-the-art industry solutions
Next, we evaluated industry state-of-the-art (SOTA) RAG solutions on CRAG public test set. We selected four RAG systems built upon SOTA LLMs and search engines, queried them with CRAG questions, collected the responses, and applied manual grading (details in Appendix A.4).
In addition, we applied traffic weights to the questions to understand the solutions in real-world use cases. The traffic weights come from a real QA use case and are generated as follows. Within each domain, we first clustered the questions into sub-domains (e.g., current game points, sports teams), and then derived the sub-domain weights from aggregated data reflective of user interactions. We applied the sub-domain weight to each CRAG question to bridge the result to reflect user experience, and reported the macro average scores across all domains (i.e., giving the same weight to all domains) in Table 6.
Table 6: Benchmarking CRAG questions with industry SOTA RAG systems. Perfect, acceptable (Acc.), hallucination (Hall.), missing (Miss.) rates, and scoreh are in percentages. The best system achieves a score of 51% and provides perfect answers for up to 63% of questions.
System Perfect Acc. Hall. Miss. Scoreh Latency (ms)
@@
ChatGPT Plus 59.8 13.1 25.2 1.9 41.2 6,195
Perplexity.ai 55.8 8.8 25.3 10.1 34.9 2,455
Traffic Copilot Pro 60.7 14.2 18.5 6.7 49.3 -
weighted Gemini Advanced 59.1 10.9 16.6 13.4 48.0 -
ChatGPT Plus 59.7 13.3 25.1 1.9 41.3 -
Perplexity.ai 55.8 9.7 25.2 9.3 35.5 -
Figure 3: SOTA systems human-eval scores (in percentage) across different dimensions.
Table 6 and Figure 3 show the overall performance of the SOTA systems and their performance across different dimensions. The evaluation results confirm our belief that the CRAG benchmark reveals interesting insights and shows room for improvement for existing RAG solutions. First, the scores from SOTA solutions achieve much better scores (highest $51\%$) compared to the straightforward solutions. However, the weighted hallucination rate ranges from 17% to 25%, so the answers are still not trustworthy. Note that the scores between the SOTA solutions and the straightforward solutions are not completely comparable, as they have different accesses to retrieval contents (Appendix A.3 and A.4.1), and the former used auto-eval, while the latter used human-eval; however, the trend is valid. Second, most difficult slices we see in the straightforward solutions remain to be difficult for SOTA solutions: real-time and fast-changing queries, and questions regarding torso and tail entities, showing the improvement needed for handling retrieval noises when the system relies on retrieval results to answer the question; as another example, we see lower scores for queries requiring multi-hop reasoning or post-processing, showing the improvement space for reasoning in question answering. Third, the third SOTA system has slightly higher weighted accuracy (perfect + acceptable) compared to the second system ($73\%$ and $70\%$), but much higher hallucination ($25.1\%$ and $16.6\%$), indicating the needs for building RAG systems to wisely answer “I don’t know" when confident answers cannot be found. Last, scores on set and false premise questions improved significantly in the SOTA solutions compared to the straightforward solutions, showing advancement in RAG systems in providing accurate and complete set answers and detecting false premises.
Finally, We observed very different latency, ranging from $2.5$s to $11.6$s, reflecting the different design options in trading off latency and quality. Note that the latency results come from interacting with the API for Perplexity.ai and the web interfaces for the other systems. See Appendix A.4.2 for additional results and how we measured latency.
@@ A.1.3 Constructing QA pairs from web contents
A.1.4 An example of retrieved web search results
@@
A.1.5 Estimated recall for retrieved webpages
Table 9: Estimated recall (in percentage) for Web and KG Questions.
Web Questions KG Questions
Question Type Est. Recall Count Est. Recall Count
Simple 94 160 60 1,045
Simple w. Condition 89 420 81 269
Set 87 326 23 186
Comparison 97 350 84 158
Aggregation 95 331 91 149
Multi-hop 91 312 83 77
Post-processing heavy 75 150 90 30
False Premise 49 376 0 70
All 84 2,425 63 1,984
@@ A.2.1 Human evaluation
We run human evaluation to score each answer with respect to the metrics defined in Section 4.1. We score each perfect, acceptable, missing, or incorrect answer with a score $s_{p}$, $s_{a}$, $s_{m}$, and $s_{in}$, respectively and define Scoreh as the score of the answer by setting $s_{p}=1$, $s_{a}=0.5$, $s_{m}=0$, and $s_{in}=-1$. We then compute the average score for all examples in the evaluation set as the score for the RAG solution.
@@ A.2.2 Automatic evaluation
In auto-eval, we merge perfect and acceptable as accurate and consider only three scores: 1 for accurate, 0 for missing, and -1 for incorrect. The resulting Scorea is calculated by the average score for all the examples in the evaluation set, and is effectively
@@
where Accuracy, Hallucination, and Missing are the percentage of accurate, incorrect, and missing answers in the test set. These score choices penalize incorrect answers, while awarding correct answers only.
We computed the accuracy and F1 for the two auto-evaluators for the accurate, incorrect, and missing examples in the public test set, respectively. Here, we considered human-eval labels as the ground truth. Table 10 shows both models attain reasonably good accuracy and F1 scores as an evaluator compared to the manual labels.
@@ A.3.1 Prompts used in straightforward RAG solutions
Vanilla LLM. PROMPT = """ You are given a Question and the time when it was asked in the Pacific Time Zone (PT), referred to as ’Query Time’. The query time is formatted as "mm/dd/yyyy, hh:mm:ss PT". Your task is to answer the question in as few words as possible.
@@
For ‘encyclopedia‘ and ‘other‘ queries, these are possible keys:
- ‘main_entity‘: extract the main entity of the query.
For ‘finance‘ queries, these are possible keys:
- ‘market_identifier‘: stock identifiers including individual company names, stock symbols.
- ‘metric‘: financial metrics that the query is asking about. This must be one of the following: ‘price‘, ‘dividend‘, ‘P/E ratio‘, ‘EPS‘, ‘marketCap‘, and ‘other‘.
- ‘datetime‘: time frame that the query asks about. When datetime is not explicitly mentioned, use ‘Query Time‘ as default.
For ‘movie‘ queries, these are possible keys:
- ‘movie_name‘: name of the movie
- ‘movie_aspect‘: if the query is about a movie, which movie aspect the query asks. This must be one of the following: ‘budget‘, ‘genres‘, ‘original_language‘, ‘original_title‘, ‘release_date‘, ‘revenue‘, ‘title‘, ‘cast‘, ‘crew‘, ‘rating‘, ‘length‘.
- ‘person‘: person name related to moves
- ‘person_aspect‘: if the query is about a person, which person aspect the query asks. This must be one of the following: ‘acted_movies‘, ‘directed_movies‘, ‘oscar_awards‘, ‘birthday‘.
- ‘year‘: if the query is about movies released in a specific year, extract the year
For ‘music‘ queries, these are possible keys:
- ‘artist_name‘: name of the artist
- ‘artist_aspect‘: if the query is about an artist, extract the aspect of the artist. This must be one of the following: ‘member‘, ‘birth place‘, ‘birth date‘, ‘lifespan‘, ‘artist work‘, ‘grammy award count‘, ‘grammy award date‘.
- ‘song_name‘: name of the song
- ‘song_aspect‘: if the query is about a song, extract the aspect of the song. This must be one of the following: ‘author‘, ‘grammy award count‘, ‘release country‘, ‘release date‘.
For ‘sports‘ queries, these are possible keys:
- ‘sport_type‘: one of ‘basketball‘, ‘soccer‘, ‘other‘
- ‘tournament‘: NBA, World Cup, Olympic.
- ‘team‘: teams that users are interested in.
- ‘datetime‘: time frame that the user is interested in. When datetime is not explicitly mentioned, use ‘Query Time‘ as default.
@@ A.3.2 Performance of straightforward RAG solutions
Table 11 summarizes the results of straightforward RAG solutions. We measured the latency on A100 (80GB) GPUs.
@@
Model Accuracy (%) Hallucination (%) Missing (%) Score (%) Latency (ms)
LLM only Llama 2 7B Chat 14.8 78.4 6.7 -63.6 289
Llama 2 70B Chat 22.3 28.7 49.0 -6.4 1,392
Llama 3 8B Instruct 23.7 33.8 42.6 -10.1 823
Llama 3 70B Instruct 32.3 28.9 38.8 3.4 1,252
GPT-4 Turbo 33.5 13.5 53.0 20.0 619
Task 1 Llama 2 7B Chat 16.4 83.1 0.5 -66.7 1,545
Llama 2 70B Chat 29.3 61.0 9.7 -31.7 5,229
Llama 3 8B Instruct 28.5 45.6 25.9 -17.1 2,152
Llama 3 70B Instruct 35.6 31.1 33.3 4.5 2,481
GPT-4 Turbo 35.9 28.2 35.9 7.7 1,270
Task 2 Llama 2 7B Chat 16.4 83.1 0.5 -66.7 1,569
Llama 2 70B Chat 29.1 61.1 9.7 -32.0 5,251
Llama 3 8B Instruct 28.6 45.5 25.9 -16.9 1,701
Llama 3 70B Instruct 37.5 29.2 33.3 8.3 2,175
GPT-4 Turbo 41.3 25.1 33.6 16.2 1,074
Task 3 Llama 2 7B Chat 16.0 83.6 0.4 -67.6 2,142
Llama 2 70B Chat 31.9 65.7 2.4 -33.7 6,149
Llama 3 8B Instruct 32.1 56.3 11.6 -24.1 4,536
Llama 3 70B Instruct 40.6 31.6 27.8 9.1 5,028
GPT-4 Turbo 43.6 30.1 26.3 13.4 1,688

### v2

@@ 5.2 State-of-the-art industry solutions
Next, we evaluated industry state-of-the-art (SOTA) RAG solutions on CRAG public test set. We selected five RAG systems built upon SOTA LLMs and search engines, queried them with CRAG questions, collected the responses, and applied manual grading (details in Appendix A.4).
In addition, we applied traffic weights to the questions to understand the solutions in real-world use cases. The traffic weights reflect the frequency of each question type, as defined in Table 2, in real QA traffic. We gave the same weights to all domains and reported the macro average across domains. This is because we have observed quite different domain-level distributions in different use cases, but have been observing similar distributions at the query-type level.
Table 6: Benchmarking CRAG questions with industry SOTA RAG systems. Perfect, acceptable (Acc.), hallucination (Hall.), missing rates (Miss.), and truthfulnessh reported by human-eval (Truthh) are in percentages. The best system achieves truthfulness of 51% and provides perfect answers for up to 63% of questions.
System Perfect Acc. Hall. Miss. Truthh Latency (ms)
@@
ChatGPT Plus 59.8 13.3 25.0 1.9 41.5 6,195
Meta SG 52.5 9.7 16.0 21.8 41.4 3,431
Perplexity.ai 55.8 8.8 25.3 10.1 34.9 4,634
Traffic Copilot Pro 70.0 9.5 14.3 6.1 60.5 -
weighted Gemini Advanced 67.1 10.0 12.7 10.2 59.3 -
ChatGPT Plus 61.8 11.4 25.7 1.3 41.8 -
Meta SG 61.0 7.1 14.1 17.8 50.5 -
Perplexity.ai 63.7 6.3 20.9 9.1 45.9 -
Figure 4: SOTA systems human-eval truthfulness scores (in percentage) across different dimensions.
Table 6 and Figure 4 show the overall performance of the SOTA systems and their performance across different dimensions. The evaluation results confirm our belief that the CRAG benchmark reveals interesting insights and shows room for improvement for existing RAG solutions. First, the results from SOTA solutions achieve much better truthfulness (highest $51\%$) compared to the straightforward solutions. However, the hallucination rate ranges from 16% to 25%, so the answers are still not trustworthy. Note that the truthfulness scores between the SOTA solutions and the straightforward solutions are not completely comparable, as they have different accesses to retrieval contents (Appendix A.3 and A.4.1), and the former used auto-eval, while the latter used human-eval; however, the trend is valid. Second, we observed very different latency, ranging from $3.4$s to $11.6$s, reflecting the different design options in trading off latency and quality; for example, Copilot Pro has the highest truthfulness, but meanwhile highest latency, whereas Meta SG [10] has mid-tier truthfulness but lowest latency. (See Appendix A.4.2 for additional results and how we measured latency.) Third, most difficult slices we see in the straightforward solutions remain to be difficult for SOTA solutions: real-time and fast-changing queries, and questions regarding torso and tail entities, showing the improvement needed for handling retrieval noises when the system relies on retrieval results to answer the question; as another example, we see lower truthfulness for queries requiring aggregation, multi-hop reasoning or post-processing, showing the improvement space for reasoning in question answering. Last, truthfulness on set and false premise questions improved significantly in the SOTA solutions compared to the straightforward solutions, showing advancement in RAG systems in providing accurate and complete set answers and detecting false premises.
@@ 6 Conclusion
Acknowledgements
We would like to thank Jinsong Yu for his advice on the data strategy and help in connecting with the partner teams. We thank Alex Boesenberg for enabling us to create web search results. We thank Sam Wexler for coordinating this project with various partners. We thank Hejia Zhang, Rares Bostan, Eryk Helenowski, Nanshu Wang, Sinong Wang, and Denis Savenkov for the discussion regarding LLM and RAG evaluation. We thank Katie Manlove for coordinating with the budget and annotation needs. We thank our annotation team for creating many great examples: David Vu, Florian Gawlitta, Rani Salomi Pitta, Lauren Gregory Mikus, Tara Welch, Nidhi Modi, Ray De Leon, Jader Ricarte, Joshua Aranzaso. We thank Apoorva Srinivas and Courtnee Parker for coordinating the annotation tasks. We would like to acknowledge the support from Jackie Leader, Mindy Abern, Heather Nolan, Ty Toledano, George Lan, Ben Edgar, Sy Choudhury, Rodrick Shepard, Katie Manlove, Mavis Hu, Jen Vinz, Taha Masood, Parkin Kent, Rebekkah Hogan, Tony Nelli, Jennifer Pak, Jonathan Torres, and Amy Lee.
Lei Chen’s work is partially supported by National Key Research and Development Program of China Grant No. 2023YFF0725100, National Science Foundation of China (NSFC) under Grant No. U22B2060, the Hong Kong RGC GRF Project 16213620, RIF Project R6020-19, AOE Project AoE/E-603/18, Theme-based project TRS T41-603/20R, CRF Project C2004-21G, Guangdong Province Science and Technology Plan Project 2023A0505030011, Hong Kong ITC ITF grants MHX/078/21 and PRP/004/22FX, Zhujiang scholar program 2021JC02X170, Microsoft Research Asia Collaborative Research Grant, HKUST-Webank joint research lab and HKUST(GZ)-Chuanglin Graph Data Joint Lab.
@@ A.1.3 Constructing QA pairs from web contents
A.1.4 Validation for the QA pairs
We conducted two phases of dataset validation with our in-house Linguist team.
Phase 1. Question and meta-label validation. After the initial round of QA pair generation, an audit session was conducted, where expert annotators reviewed the question template, questions, and meta-labels (domain, question type, etc), applying edits as necessary with 2x human review (agreement rate $90\%+$). All problematic questions (e.g., a wrong false-premise question) were revised, and all conflicting labels were resolved by a third more experienced auditor.
Phase 2. Answer validation. To ensure the answers in the benchmark are correct, we further conducted an auditing process for all the answers. For the web questions, an annotation team reviewed each question and conducted an extensive search to make sure the answer is factually correct and includes comprehensive information (such as for set questions) with 2x human review (agreement rate $90\%+$). A third more experienced auditor then reviewed all conflicting answers and provided a resolution. For the KG questions, a team of five engineers carefully checked the questions and queried the mock APIs manually to validate the answers. This step resulted in a $5\%$ answer correction.
In both phases, we paid special attention to examples where the straightforward solutions output different answers from the ground truth answers and asked the auditing team to double-check those examples. This step yielded an additional $2\%$ answer updates.
A.1.5 An example of retrieved web search results
@@ A.1.6 The mock data and mock APIs
Table 9: Mock APIs and Descriptions.
APIs Descriptions
open_search_entity_by_name(query: str) -> dict Search for entities by name in the Open domain.
open_get_entity(entity: str) -> dict Retrieve detailed information about an entity in the Open domain.
movie_get_person_info(person_name: str) -> dict Get information about a person related to movies.
movie_get_movie_info(movie_name: str) -> dict Get information about a movie.
movie_get_year_info(year: str) -> dict Get information about movies released in a specific year.
movie_get_movie_info_by_id(movie_id: int) -> dict Get movie information by its unique ID.
movie_get_person_info_by_id(person_id: int) -> dict Get person information by their unique ID.
finance_get_company_name(query: str) -> dict Search for company names in the finance domain.
finance_get_ticker_by_name(query: str) -> dict Retrieve the ticker symbol for a given company name.
finance_get_price_history(ticker_name: str) -> dict Get the price history for a given ticker symbol.
finance_get_detailed_price_history(ticker_name: str) -> dict Get detailed price history for a ticker symbol.
finance_get_dividends_history(ticker_name: str) -> dict Get dividend history for a ticker symbol.
finance_get_market_capitalization(ticker_name: str) -> dict Retrieve market capitalization for a ticker symbol.
finance_get_eps(ticker_name: str) -> dict Get earnings per share (EPS) for a ticker symbol.
finance_get_pe_ratio(ticker_name: str) -> dict Get the price-to-earnings (PE) ratio for a ticker symbol.
finance_get_info(ticker_name: str) -> dict Get financial information for a ticker symbol.
music_search_artist_entity_by_name(artist_name: str) -> dict Search for music artists by name.
music_search_song_entity_by_name(song_name: str) -> dict Search for songs by name.
music_get_billboard_rank_date(rank: int, date: str = None) -> dict Get Billboard ranking for a specific rank and date.
music_get_billboard_attributes(date: str, attribute: str, song_name: str) -> dict Get attributes of a song from Billboard rankings.
music_grammy_get_best_artist_by_year(year: int) -> dict Get the Grammy Best New Artist for a specific year.
music_grammy_get_award_count_by_artist(artist_name: str) -> dict Get the total Grammy awards won by an artist.
music_grammy_get_award_count_by_song(song_name: str) -> dict Get the total Grammy awards won by a song.
music_grammy_get_best_song_by_year(year: int) -> dict Get the Grammy Song of the Year for a specific year.
music_grammy_get_award_date_by_artist(artist_name: str) -> dict Get the years an artist won a Grammy award.
music_grammy_get_best_album_by_year(year: int) -> dict Get the Grammy Album of the Year for a specific year.
music_grammy_get_all_awarded_artists() -> dict Get all artists awarded the Grammy Best New Artist.
music_get_artist_birth_place(artist_name: str) -> dict Get the birthplace of an artist.
music_get_artist_birth_date(artist_name: str) -> dict Get the birth date of an artist.
music_get_members(band_name: str) -> dict Get the member list of a band.
music_get_lifespan(artist_name: str) -> dict Get the lifespan of an artist.
music_get_song_author(song_name: str) -> dict Get the author of a song.
music_get_song_release_country(song_name: str) -> dict Get the release country of a song.
music_get_song_release_date(song_name: str) -> dict Get the release date of a song.
music_get_artist_all_works(artist_name: str) -> dict Get all works by an artist.
sports_soccer_get_games_on_date(team_name: str, date: str) -> dict Get soccer games on a specific date.
sports_nba_get_games_on_date(team_name: str, date: str) -> dict Get NBA games on a specific date.
sports_nba_get_play_by_play_data_by_game_ids(game_ids: List[str]) -> dict Get NBA play by play data for a set of game ids.
@@ A.2.1 Human evaluation
We run human evaluation to score each answer with respect to the metrics defined in Section 4.1. We score each perfect, acceptable, missing, or incorrect answer with a score $s_{p}$, $s_{a}$, $s_{m}$, and $s_{in}$, respectively and define truthfulnessh as the score of the answer by setting $s_{p}=1$, $s_{a}=0.5$, $s_{m}=0$, and $s_{in}=-1$. We then compute the average truthfulness for all examples in the evaluation set as the truthfulness score for the RAG solution.
Human-eval instructions. The human-eval instructions are as follows.
Given a query, a query day and time at which the query was made, the chatbot’s response, grade the accuracy for the response according to the criteria below:
Using an external search engine, please evaluate the factual accuracy of the response based on the grading rubric below. An accurate answer should be factually correct and provide useful information to answer the user’s question.
  • Accuracy = 0 (Missing). It covers the following situations. There is no response. There is a failure to provide a response to the request (e.g. “I’m sorry ……, I can’t do …”) due to inability. E.g., Query: “Latest news about the Nobel prize today.” Response: “I can’t find specific information regarding the Nobel prize …” The response fails to answer and asks follow-up questions.
  • Accuracy = 1 (Incorrect). It covers the following situations. The answer is unintelligible. The answer is poorly formed. The answer contains a major hallucination (e.g. wrong date, wrong numbers, or other significant factual errors). The answer is irrelevant to the user’s request. Answers in a category, location, or time window that is significantly different from the user’s request, if any. There is a significant structural/formatting error. The response is otherwise a total structural/functional failure and does not contain sufficient well-formed content that can be used to determine accuracy
  • Accuracy = 2 (Acceptable). It covers the following situations. The answer is acceptably correct and relevant to the user’s request, but may miss some information, i.e. accurate but not complete, or mostly accurate with minor issues. The answer may contain some minor hallucination that doesn’t significantly alter the overall meaning. “Minor hallucination” means the answer addressed the user’s question but might be off on some additional details. The rule of thumb is to see if the answer serves the purpose of the user’s question, and whether the hallucination could mislead the users on what they were asking.
  • Accuracy = 3 (Accurate). The answer is factually correct, contains all the relevant information requested, responds appropriately to the query, and does not contain any hallucination.
We requested two human graders to grade each question (agreement rate $94\%$), and when there is a conflict, a third more experienced grader will resolve it.
@@ A.2.2 Automatic evaluation
In auto-eval, we merge perfect and acceptable as accurate and consider only three scores: 1 for accurate, 0 for missing, and -1 for incorrect. The truthfulnessa is calculated by the average truthfulness for all the examples in the evaluation set, and is effectively
@@
where accuracy, hallucination, and missing are the percentage of accurate, incorrect, and missing answers in the test set. These score choices penalize incorrect answers, award correct, and assign a value of 0 to missing answers.
Auto-evaluators. We computed the accuracy and F1 for the two auto-evaluators for the accurate, incorrect, and missing examples in the public test set, respectively. Here, we considered human-eval labels as the ground truth. Table 10 shows both models attain reasonable accuracy and F1 scores as an evaluator compared to human evaluation.
@@
Auto-eval prompt. The prompt we used in the auto-eval is similar to the following. We did not release the exact prompt used in the challenge to avoid prompt attack.
PROMPT = """# Task: You are given a Question, a model Prediction, and a list of Ground Truth answers, judge whether the model Prediction matches any answer from the list of Ground Truth answers. Follow the instructions step by step to make a judgement.
1. If the model prediction matches any provided answers from the Ground Truth Answer list, "Accuracy" should be "True"; otherwise, "Accuracy" should be "False".
2. If the model prediction says that it couldn’t answer the question or it doesn’t have enough information, "Accuracy" should always be "False".
3. If the Ground Truth is "invalid question", "Accuracy" is "True" only if the model prediction is exactly "invalid question".
# Output:
Respond with only a single JSON string with an "Accuracy" field which is "True" or "False".
# Examples:
Question: how many seconds is 3 minutes 15 seconds?
Ground truth: ["195 seconds"]
Prediction: 3 minutes 15 seconds is 195 seconds.
Accuracy: True
Question: Who authored The Taming of the Shrew (published in 2002)?
Ground truth: ["William Shakespeare", "Roma Gill"]
Prediction: The author to The Taming of the Shrew is Roma Shakespeare.
Accuracy: False
Question: Who played Sheldon in Big Bang Theory?
Ground truth: ["Jim Parsons", "Iain Armitage"]
Prediction: I am sorry I don’t know.
Accuracy: False
@@ A.3.1 Prompts used in straightforward RAG solutions
Vanilla LLM. PROMPT = """ You are given a Question and the time when it was asked in the Pacific Time Zone (PT), referred to as "Query Time". The query time is formatted as "mm/dd/yyyy, hh:mm:ss PT". Your task is to answer the question in as few words as possible.
@@
For ‘encyclopedia’ and ‘other’ queries, these are possible keys:
- ‘main_entity’: extract the main entity of the query.
For ‘finance’ queries, these are possible keys:
- ‘market_identifier’: stock identifiers including individual company names, stock symbols.
- ‘metric’: financial metrics that the query is asking about. This must be one of the following: ‘price’, ‘dividend’, ‘P/E ratio’, ‘EPS’, ‘marketCap’, and ‘other’.
- ‘datetime’: time frame that the query asks about. When datetime is not explicitly mentioned, use ‘Query Time’ as default.
For ‘movie’ queries, these are possible keys:
- ‘movie_name’: name of the movie
- ‘movie_aspect’: if the query is about a movie, which movie aspect the query asks. This must be one of the following: ‘budget’, ‘genres’, ‘original_language’, ‘original_title’, ‘release_date’, ‘revenue’, ‘title’, ‘cast’, ‘crew’, ‘rating’, ‘length’.
- ‘person’: person name related to moves
- ‘person_aspect’: if the query is about a person, which person aspect the query asks. This must be one of the following: ‘acted_movies’, ‘directed_movies’, ‘oscar_awards’, ‘birthday’.
- ‘year’: if the query is about movies released in a specific year, extract the year
For ‘music’ queries, these are possible keys:
- ‘artist_name’: name of the artist
- ‘artist_aspect’: if the query is about an artist, extract the aspect of the artist. This must be one of the following: ‘member’, ‘birth place’, ‘birth date’, ‘lifespan’, ‘artist work’, ‘grammy award count’, ‘grammy award date’.
- ‘song_name’: name of the song
- ‘song_aspect’: if the query is about a song, extract the aspect of the song. This must be one of the following: ‘author’, ‘grammy award count’, ‘release country’, ‘release date’.
For ‘sports’ queries, these are possible keys:
- ‘sport_type’: one of ‘basketball‘, ‘soccer‘, ‘other‘
- ‘tournament’: NBA, World Cup, Olympic.
- ‘team’: teams that users are interested in.
- ‘datetime’: time frame that the user is interested in. When datetime is not explicitly mentioned, use ‘Query Time’ as default.
@@ A.3.2 Performance of straightforward RAG solutions
Table 11 summarizes the results of straightforward RAG solutions.
@@
Model Accuracy (%) Hallucination (%) Missing (%) Truthfulness (%)
LLM only Llama 2 7B Chat 14.8 78.4 6.7 -63.6
Llama 2 70B Chat 22.3 28.7 49.0 -6.4
Llama 3 8B Instruct 23.7 33.8 42.6 -10.1
Llama 3 70B Instruct 32.3 28.9 38.8 3.4
Falcon 40B 10.8 41.9 47.3 -31.1
FLAN-T5-XXL 11B 9.4 8.7 81.9 0.7
Mixtral-8x7B-Instruct-v0.1 20.8 27.0 52.1 -6.2
GPT-4 Turbo 33.5 13.5 53.0 20.0
Task 1 Llama 2 7B Chat 16.4 83.1 0.5 -66.7
Llama 2 70B Chat 29.3 61.0 9.7 -31.7
Llama 3 8B Instruct 28.5 45.6 25.9 -17.1
Llama 3 70B Instruct 35.6 31.1 33.3 4.5
Falcon 40B 21.9 55.5 22.5 -33.6
FLAN-T5-XXL 11B 27.5 36.5 36.0 -9.0
Mixtral-8x7B-Instruct-v0.1 33.6 44.4 22.0 -10.8
GPT-4 Turbo 35.9 28.2 35.9 7.7
Task 2 Llama 2 7B Chat 16.4 83.1 0.5 -66.7
Llama 2 70B Chat 29.1 61.1 9.7 -32.0
Llama 3 8B Instruct 28.6 45.5 25.9 -16.9
Llama 3 70B Instruct 37.5 29.2 33.3 8.3
Falcon 40B 21.9 55.4 22.7 -33.5
FLAN-T5-XXL 11B 27.4 36.6 36.0 -9.2
Mixtral-8x7B-Instruct-v0.1 33.4 44.6 22.0 -11.2
GPT-4 Turbo 41.3 25.1 33.6 16.2
Task 3 Llama 2 7B Chat 16.0 83.6 0.4 -67.6
Llama 2 70B Chat 31.9 65.7 2.4 -33.7
Llama 3 8B Instruct 32.1 56.3 11.6 -24.1
Llama 3 70B Instruct 40.6 31.6 27.8 9.1
Falcon 40B 22.0 56.6 21.3 -34.6
FLAN-T5-XXL 11B 27.8 37.1 35.1 -9.3
Mixtral-8x7B-Instruct-v0.1 33.5 44.1 22.4 -10.6
GPT-4 Turbo 43.6 30.1 26.3 13.4